OpenStudy (anonymous):

Can someone help me with this problem: (x+3) (x^2-2x+1) Can you possibly explain it to me as well

jimthompson5910 (jim_thompson5910):

Use the idea that (A+B)(C+D+E) = A(C+D+E) + B(C+D+E) and simplify to get... (x+3)(x^2-2x+1) x(x^2-2x+1)+3(x^2-2x+1) x^3-2x^2+x+3x^2-6x+3 x^3+x^2-5x+3 So the expression expands and simplifies to x^3+x^2-5x+3

OpenStudy (anonymous):

lead me through the process, please

jimthompson5910 (jim_thompson5910):

Let me know if that helps. If not, I can try another way.

OpenStudy (anonymous):

can you explain it please

jimthompson5910 (jim_thompson5910):

In general, (A+B)(C+D+E) = A(C+D+E) + B(C+D+E)

jimthompson5910 (jim_thompson5910):

so use this to transform (x+3)(x^2-2x+1) into x(x^2-2x+1)+3(x^2-2x+1)

jimthompson5910 (jim_thompson5910):

it's like a form of distribution

jimthompson5910 (jim_thompson5910):

next, distribute each group so distribute x to x^2-2x+1 to get x(x^2-2x+1) = x^3 - 2x^2 + x -------------------------------- then distribute 3 to x^2-2x+1 to get 3(x^2-2x+1) = 3x^2 - 6x + 3

jimthompson5910 (jim_thompson5910):

so x(x^2-2x+1)+3(x^2-2x+1) then becomes x^3 - 2x^2 + x + 3x^2 - 6x + 3 from here, all you need to do is combine like terms so add -2x^2 and 3x^2 to get x^2 and add x and -6x to get -5x giving you the final answer of x^3 + x^2 - 5x + 3

jimthompson5910 (jim_thompson5910):

Let me know if this makes sense or not.

OpenStudy (anonymous):

will do, thanks